# ETH-monotonicity and the black hole singularity

### 11 Mar 2026 v3

@@ Abstract
ETH-monotonicity is a property of quantum chaotic many-body systems which reinforces the Kelvin statement of the second law of thermodynamics over and above the universal entropic contribution. We show that higher-dimensional holographic conformal field theories possess ETH-monotonicity. Smaller black hole microstates have stronger contribution from ETH-monotonicity to the second law of thermodynamics. ETH-monotonicity also dominates quantum fluctuations in small black hole microstates in higher dimensions. We find that the relative extra gain in energy measures the curvature at the black hole horizon of small black holes. In the smallest size limit, the black hole curvature singularity is constructed of microstates for which ETH-monotonicity starts competing with the entropic factor. We expect that ETH-monotonicity will persist even in the ultimate quantum theory of gravity. Because it is a property of many-body quantum chaotic systems which becomes more prominent with decreasing system size, unlike other physical properties which are usually more predictable with increasing system size. Two-dimensional holographic conformal field theory does not possess all features of ETH-monotonicity which is in agreement with the absence of curvature singularity in the BTZ black hole.
@@ 1 Introduction
It would not be an overstatement to state that the black hole singularity (and, in general, other spacetime singularities) has been the greatest mystery in physics. This singularity is found abundantly in physics. So, we expect that its physics underpinning should be robust and universal. Black holes are also known to be highly chaotic Sekino:2008he ; Maldacena:2015waa . In this paper, we provide a promising lead towards resolving this mystery using a newly discovered property of quantum chaotic many-body systems. We show that this property arises from strong curvature at the black hole horizon.
@@
Quantum chaos and the related topic of thermalization in closed quantum systems have been subjects of great interest for the last few decades Polkovnikov:2010yn ; DAlessio:2015qtq ; Gogolin:2016hwy . It is expected that closed quantum systems thermalize in the long time limit. This explains the validity of quantum statistical mechanics in nature. In this regard, eigenstate thermalization hypothesis (ETH) PhysRevA.43.2046 ; Srednicki:1994mfb explains thermalization after unitary time evolution. So, ETH has been a subject of intense study in different set-ups, mostly in finite lattice systems using exact diagonalization DAlessio:2015qtq .
ETH is concerned with observables which can be easily measured or calculated. We say that a system is thermalized if these observables have expectation values equal to the thermal expectation values. Consider a non-fermionic hermitian operator $\mathcal{O}$ which corresponds to one such observable. ETH ansatz of the matrix elements $\mathcal{O}_{mn}$ in energy eigenstates is DAlessio:2015qtq
@@
$\bar{E}=(E_{m}+E_{n})/2,\quad\omega=E_{m}-E_{n},\quad f(\bar{E},\omega)=f(\bar{E},-\omega)\geq 0\$
where $|m\rangle$ and $|n\rangle$ are energy eigenstates with energies $E_{m}$ and $E_{n}$ respectively. $S(\bar{E})$ is the entropy at energy $\bar{E}$. $\mathcal{O}(\bar{E})$ and $f(\bar{E},\omega)$ are smooth functions of their arguments. So, $\mathcal{O}(\bar{E})$ is equal to the expectation value of $\mathcal{O}$ in the microcanonical ensemble centered at energy $\bar{E}$. $R_{mn}$ are pseudo-random variables with zero mean and unit variance, and $R_{nm}=R^{*}_{mn}$. At large $|\omega|$, the $f$-function falls at least as fast as $e^{-|\beta\omega|/4}$, where $\beta$ is the effective inverse temperature Srednicki_1999 ; DAlessio:2015qtq ; Dymarsky:2018ccu ; Murthy:2019fgs . The second fluctuation term in (1.1) exists even for the diagonal elements. The ETH statement for the diagonal elements without the fluctuations is equivalent to Berry’s conjecture Berry_1977 . It implies that eigenstate expectation values are equal to the microcanonical averages in quantum chaotic systems DAlessio:2015qtq . The most important difference between ETH and random matrix theory (RMT) is the non-trivial $f$-function in ETH. In RMT, it is a constant function.
The $\omega$-dependence of the $f$-function is related with fluctuation and the response function of the operator $\mathcal{O}$ Srednicki_1999 . So, the $\omega$-dependence has been studied for many systems DAlessio:2015qtq . Recently, the $\bar{E}$-dependence was studied in lattice models in ETHmono ; sorokhaibam2025ethmono2D . It is also closely related with the second law of thermodynamics. We found an interesting set of constraints on the $\bar{E}$-dependence of the $f$-function which we called ETH-monotonicity. In this work, we will explore ETH-monotonicity in holographic CFTs. We find a concrete connection between ETH-monotonicity and high-curvature at the black hole horizon.
@@
    We will study $\bar{E}\to 0$ behaviour. We are more interested in the behaviour of the $f$-function as $\bar{E}$ decreases. We are still working with many-body excited state $\bar{E}\sim N$, where $N$ is the number of degrees of freedom of the holographic CFT. In the AdS side, it means that we are still working with classical gravity. We are not considering stringy or quantum gravity corrections.
The outline of the rest of the paper is as follows. We review the necessary basics of AdS/CFT correspondence in section 2. ETH-monotonicity is defined and studied in section 3. Section 4 contains summary of our results and some discussions. Quantum fluctuations in small black holes are studied in section 5. Section 6 contains our main results. We show that higher-d holographic CFT possesses ETH-monotonicity. The special case of 2-d holographic CFT is studied in section 7. Section 8 consists of conclusions from this work. ETH-monotonicity in microcanonical ensemble is worked out in Appendix A. Further numerical results including the case of 3-d holographic CFT are presented in Appendix B.
@@ 2 Black holes and AdS/CFT correspondence
where $\Gamma$ is the Gamma function and $G_{N}$ is the Newton’s constant. The horizon radius is the largest zero of $f(r)$. In this work, we will not consider black holes with charges or angular momentum. The Hawking temperature and the Bekenstein-Hawking entropy of the black hole are given in terms of $r_{h}$ by
@@
In the rest of the paper, we will set the AdS radius $R=1$. For $D\geq 4$, $T$ has the global minimum at $r_{h,min}=\sqrt{(d-2)/d}$. $T$ increases as $r_{h}$ decreases for small black holes $r_{h}<r_{h,min}$. So, there are three possible phases at any given temperature - thermal AdS, small black hole, and large black hole. The small black holes have negative specific heat like black holes in asymptotically flat spacetime and they are always unstable in the canonical ensemble. In higher dimensions $D\geq 4$, canonical ensemble is ill-defined for small black holes due to the superlinear entropy growth with increasing energy,
$S(\bar{E})\sim\bar{E}^{(d-1)/(d-2)},\qquad d=D-1.$ (2.4)
@@
The stable phases are the AdS thermal gas and the large black hole phase which are separated by a first order Hawking-Page transition Hawking:1982dh . The critical temperature is given by $r_{h}=1$ which is slightly away from $r_{h,min}$. But the small black holes are stable in the microcanonical ensemble. As such closed systems, they are also stable from perturbations Regge:1957td . Dynamically, their lifetimes are very long even in the canonical ensemble Horowitz:1999uv . Small charged black holes are unstable from perturbations Cardoso:2004hs ; Cardoso:2006wa ; Basu:2010uz , the end-points are still small hairy black holes. For $D=3$, the Hawking-Page transition is still between the thermal AdS phase and the black hole phase at the critical point $r_{h}=1$. The BTZ black hole Banados:1992wn in $D=3$ always have a positive specific heat. We are not interested in black strings and black branes so we are not concerned with Gregory-Laflamme instability Gregory:1993vy .
@@
We would still be working with small black holes. In terms of thermodynamics, we will work in microcanonical ensemble. Small black holes are of particular interest for the present work because small black holes have high curvature at the horizon. Note that in the other limit $r_{h}\to\infty$, canonical ensembles are well defined in any dimensions. Moreover, equivalence of ensembles holds true.
@@
By AdS/CFT correspondence, black holes in the $(d+1)$-D spacetime corresponds to a thermal state on the $d$-dimensional ($d$-d) holographic conformal field theory (CFT) at the AdS boundary. The temperature of the CFT is equal to the Hawking temperature of the black hole. The total energy of the CFT system $\bar{E}_{\beta}$ is equal to the mass of the black hole $M=\mu/\tilde{G}_{N}$ Balasubramanian:1999re . From the definition of $\bar{E}=(E_{n}+E_{m})/2$ and $\omega=E_{m}-E_{n}$, $|\omega|\leq 2\bar{E}$. But we do not have access to the discrete energy spectrum of quantized gravity. $\bar{E}$ (energy of heavy operators) is enhanced by a $1/G_{N}$ factor as compared to $\omega$ (energy of light operators) when measured in the unit of $1/R$. Hence, we can effectively take $\omega\in(-\infty,\infty)$. The use of these two different energy units were implicit in other works, for example, see Horowitz:1999jd ; Jokela:2015sza .
@@
In this work, we will consider perturbation with a light operator on the CFT. We are working with linear response theory so the strength of the source is chosen to be small. It has been shown that global perturbation with a light operator can form black holes Bhattacharyya:2009uu . One can use geodesic approximation Balasubramanian:1999zv in case of perturbation with heavy operators. The technical challenge is that there are no real geodesic which connects two timelike boundary points Balasubramanian:2012tu .
@@
We use the spectral function $A(\omega)$ of a scalar operator in the holographic CFT to calculate $f(\bar{E})$ of the scalar operator. In a thermal state with inverse temperature $\beta$ and energy expectation value $\bar{E}_{\beta}$,
$A(\omega)=2\sinh(\beta\omega/2)f(\bar{E}_{\beta},\omega)^{2},\qquad A(\omega)=-2\,\text{Im}\;\tilde{G}_{R}(\omega)$ (2.5)
$G_{R}(t,t^{\prime})=-i\theta(t-t^{\prime})\langle[\mathcal{O}(t),\mathcal{O}(t^{\prime})]\rangle$ (2.6)
@@
The retarded Green function $\tilde{G}_{R}(\omega)$ is calculated using the Son-Starinets method Son:2002sd . We consider a minimally-coupled massive scalar field $\phi$ of mass squared $m^{2}$ in the $d+1$-D bulk spacetime. The equation of motion in terms of the Fourier modes of $t$ and the spherical harmonics of the (d1)(d-1)-sphere is
@@
$\qquad\tilde{G}_{R,\Delta+}(\omega,l)=-\frac{\mathcal{B}(\omega,l)}{\mathcal{A}(\omega,l)}$ (2.12)
If $\mathcal{O}_{\Delta-}$ is well defined, then its retarded Green function is given by
$\tilde{G}_{R,\Delta-}(\omega,l)=\frac{\mathcal{A}(\omega,l)}{\mathcal{B}(\omega,l)}$ (2.13)
The $f$-function can be calculated using (2.5), where we have suppressed the $l$-dependence. We are looking for $\mu$, $\beta$, $T$ or $r_{h}$ dependence. While in other works, $\omega$-dependence is the priority. So it is important that (2.12) is nicely normalized $\text{lim}_{t\to 0+}G_{R}(t)=-1/t^{2\Delta}$, while using the $r$ coordinate Dodelson:2022yvn . It should be noted that a different coordinate system was used in Son:2002sd . For only the spectral function, we can also perform the calculation near the horizon due to flux conservation. In the rest of the paper, we will consider an operator $\mathcal{O}_{\Delta}$ of conformal weight $\Delta$ which can be either $\Delta_{+}$ or $\Delta_{-}$.
We will write the $f$-function as $f(\bar{E})$ when $\omega$ is fixed and the $\omega$-dependence is not important, similarly, $f(\omega)$ when the $\bar{E}$ is fixed. We will be varying $\mu$ so we are really working with $f(\bar{E})=f(\mu/\tilde{G}_{N})=\hat{f}(\mu)$. With a slight abuse of notation, we will still use $f(\mu)$ instead of $\hat{f}(\mu)$. The quantity that appears in physical processes is $f^{\prime}(\bar{E})/f(\bar{E})$. So we will use the relation

### 25 Mar 2026 v4

@@ Abstract
We study the enveloping function of the fluctuation term in eigenstate thermalization hypothesis (ETH) statement for holographic conformal field theories. We use this function to identify and examine black hole microstates. We set down a set of desirable criteria for this function called ETH-monotonicity. It reinforces the Kelvin statement of the second law of thermodynamics over and above the universal entropic contribution. We show that higher-dimensional holographic conformal field theories possess ETH-monotonicity. Stronger contribution from ETH-monotonicity to the second law of thermodynamics is observed in smaller black hole microstates. It dominates other quantum fluctuations. It also measures the curvature at the horizon of the small black holes. In the smallest size limit, the black hole curvature singularity is constructed of microstates for which ETH-monotonicity starts competing with the entropic contribution. We expect that ETH-monotonicity will persist even in the ultimate quantum theory of gravity. Because it is a property of many-body quantum chaotic systems which becomes more prominent with decreasing system size, unlike other physical properties which are usually more predictable with increasing system size. Two-dimensional holographic conformal field theory does not possess all features of ETH-monotonicity which is in agreement with the absence of curvature singularity in the BTZ black hole.
@@ 1 Introduction
It would not be an overstatement to state that the black hole singularity (and, in general, other spacetime singularities) has been the greatest mystery in physics. This singularity is found abundantly in physics. So, we expect that its physics underpinning should be robust and universal. Black holes are also known to be highly chaotic Sekino:2008he ; Maldacena:2015waa . In this work, we provide a promising lead towards resolving this mystery using a newly discovered property of quantum chaotic many-body systems. We show that this property arises from strong curvature at the black hole horizon.
@@
Quantum chaos and the related topic of thermalization in closed quantum systems have been subjects of great interest for the last few decades Polkovnikov:2010yn ; DAlessio:2015qtq ; Gogolin:2016hwy . It is expected that closed chaotic quantum systems thermalize in the long time limit. This explains the validity of quantum statistical mechanics in nature. In this regard, eigenstate thermalization hypothesis (ETH) PhysRevA.43.2046 ; Srednicki:1994mfb explains thermalization after unitary time evolution. So, ETH has been a subject of intense study in different set-ups, mostly in finite lattice systems using exact diagonalization DAlessio:2015qtq .
ETH is concerned with few-body observables which can be easily measured or calculated. We say that a system is thermalized if these observables have expectation values equal to the thermal expectation values. Consider a non-fermionic hermitian operator $\mathcal{O}$ which corresponds to one such observable. ETH ansatz of the matrix elements $\mathcal{O}_{mn}$ in energy eigenstates is DAlessio:2015qtq  [^1] We are assuming that the system obeys time-reversal invariance. Otherwise, $f(\bar{E},\omega)=f^{*}(\bar{E},-\omega)$ and $R_{nm}=R^{*}_{mn}$. It does not affect our calculations because all our quantities will be in terms of $|f(\bar{E},\omega)|$.
@@
$\bar{E}=(E_{m}+E_{n})/2,\quad\omega=E_{m}-E_{n},\quad f(\bar{E},\omega)=f(\bar{E},-\omega)\geq 0,\quad R_{nm}=R_{mn},\$
where $|m\rangle$ and $|n\rangle$ are energy eigenstates with energies $E_{m}$ and $E_{n}$ respectively. $S(\bar{E})$ is the entropy at energy $\bar{E}$. $\mathcal{O}(\bar{E})$ and $f(\bar{E},\omega)$ are smooth functions of their arguments. So, $\mathcal{O}(\bar{E})$ is equal to the expectation value of $\mathcal{O}$ in the microcanonical ensemble centered at energy $\bar{E}$. $R_{mn}$ are pseudo-random variables with zero mean and unit variance. At large $|\omega|$, the $f$-function falls at least as fast as $e^{-|\beta\omega|/4}$, where $\beta$ is the effective inverse temperature Srednicki_1999 ; DAlessio:2015qtq ; Dymarsky:2018ccu ; Murthy:2019fgs . The second fluctuation term in (1.1) exists even for the diagonal elements. The ETH statement for the diagonal elements without the fluctuations is equivalent to Berry’s conjecture Berry_1977 . It implies that eigenstate expectation values are equal to the microcanonical averages in quantum chaotic systems DAlessio:2015qtq .
The $\omega$-dependence of the enveloping function $f(\bar{E},\omega)$ controls the response function of the operator $\mathcal{O}$ Srednicki_1999 . So, the $\omega$-dependence has been studied for many systems DAlessio:2015qtq . We concentrate on the $\bar{E}$-dependence in this work. The main technical advance made in this work is that we demonstrate the use of the $\bar{E}$-dependence of the $f$-function to identify and examine black hole microstates. We use this technique to identify small black hole microstates. Then we identify a quantity of the holographic CFT which measures the strong curvature at the horizon of these small black holes. This quantity arises from inequivalence of ensembles so it increases with decreasing size of the black hole, thereby measuring the increasing curvature at the horizon of the black hole. We will write the $f$-function as $f(\bar{E})$ when $\omega$ is fixed and the $\omega$-dependence is not important. We calculate $f(\bar{E},\omega)$ from the spectral function. The spectral function is the imaginary part of the retarded Green function. The retarded Green function is calculated using the well-known Son-Starinets prescription Son:2002sd .
Recently, the $\bar{E}$-dependence was studied in lattice models in ETHmono ; ETHmono2D . It is also closely related with the second law of thermodynamics. We found an interesting set of constraints on the $\bar{E}$-dependence of the $f$-function which we called ETH-monotonicity. In this work, we will explore ETH-monotonicity in holographic CFTs. We find a concrete connection between ETH-monotonicity and high-curvature at the black hole horizon. The present work is largely independent even though we got the inspiration from the above previous works. The lattice models considered in those works have various energy scales. The scaling of $f(\bar{E})$ with varying volume of the lattice model was studied. In the present work, we are studying conformal field theory so we look for scaling of $f(\bar{E})$ with varying $\bar{E}$.
@@
    We will study $\bar{E}\to 0$ behaviour. This notation is for conciseness. We are more interested in the behaviour of the $f$-function as $\bar{E}$ decreases. We are still working with many-body excited states. In the AdS side, we are still working with the small black hole microstates in the Hilbert space of string theory Aharony_2000 . We are not considering stringy or quantum gravity corrections.
It is well known that certain geodesics can explore the black hole singularity Fidkowski:2003nf ; Festuccia:2005pi ; Rodriguez-Gomez:2021pfh ; Horowitz:2023ury ; Ceplak:2024bja ; Afkhami-Jeddi:2025wra ; Dodelson:2025jff . But the exact nature of the singularity has been as elusive as ever. The study of these “bouncing” geodesics is related to the $\omega$-dependence of $f(\bar{E},\omega)$, while we are concerned with the $\bar{E}$-dependence. Most of these works also considered the simplified black brane geometry. Moreover, it has been found that a certain black brane geometry with curvature singularity does not have any bouncing geodesic Grozdanov:2026cut . Other related works are Grinberg:2020fdj ; David:2022nfn ; Singhi:2024sdr ; Jia:2025jbi . We do not consider black branes because all black branes are related by a simple coordinate rescaling and all of them are already in the thermodynamic limit.
The outline of the rest of the paper is as follows. We review the necessary basics of AdS/CFT correspondence in section 2. ETH-monotonicity is defined and studied in section 3. Section 4 contains summary of our results and some discussions. Quantum fluctuations in small black holes are studied in section 5. Section 6 contains our main results. We show that higher dimensional holographic CFTs possess ETH-monotonicity. The special case of two dimensional holographic CFT is studied in section 7. Section 8 consists of conclusions from this work. Microcanonical ensemble is studied in detail in Appendix A. Further numerical results including the case of 3-d holographic CFT are presented in Appendix B.
@@ 2 Black holes and AdS/CFT correspondence
where $\Gamma$ is the Gamma function and $G_{N}$ is the Newton’s constant. The horizon radius is the largest zero of $g(r)$. In this work, we will not consider black holes with charges or angular momentum. The Hawking temperature and the Bekenstein-Hawking entropy of the black hole are given in terms of $r_{h}$ by
@@
In the rest of the paper, we will set the AdS radius $R=1$. For $D\geq 4$, $T$ has the global minimum at $r_{h,min}=\sqrt{(d-2)/d}$. $T$ increases as $r_{h}$ decreases for small black holes $r_{h}<r_{h,min}$. So, there are three possible phases at any given temperature - thermal AdS, small black hole, and large black hole. The small black holes have negative specific heat like black holes in asymptotically flat spacetime and they are always unstable in the canonical ensemble. In other words, canonical ensemble is ill-defined for small black holes in $D\geq 4$ spacetime due to the superlinear entropy growth with increasing energy,
$S(\bar{E})\sim(\bar{E}l_{P})^{(d-1)/(d-2)},\qquad d=D-1.$ (2.4)
@@
$l_{P}=G_{N}^{1/(d-1)}$ is the Planck length scale. The stable phases are the AdS thermal gas and the large black hole phase which are separated by a first order Hawking-Page transition Hawking:1982dh . The critical temperature is given by $r_{h}=1$ which is slightly away from $r_{h,min}$. But the small black holes are stable in the microcanonical ensemble. As such closed systems, they are also stable from perturbations Regge:1957td . Dynamically, their lifetimes are very long even in the canonical ensemble Horowitz:1999uv . Small charged black holes are unstable from perturbations Cardoso:2004hs ; Cardoso:2006wa ; Basu:2010uz , the end-points are still small hairy black holes. For $D=3$, the Hawking-Page transition is still between the thermal AdS phase and the black hole phase at the critical point $r_{h}=1$. The BTZ black hole Banados:1992wn in $D=3$ always have a positive specific heat. We are not interested in black strings and black branes so we are not concerned with Gregory-Laflamme instability Gregory:1993vy .
@@
We would still be working with small black holes. In terms of thermodynamics, we will work in microcanonical ensemble. Small black holes are of particular interest for the present work because small black holes have high curvature at the horizon. Note that in the other limit $r_{h}\to\infty$, canonical ensembles are well defined in any dimensions.
@@
By AdS/CFT correspondence, black holes in the $(d+1)$-D spacetime corresponds to a thermal state on the $d$-dimensional ($d$-d) holographic conformal field theory (CFT) at the AdS boundary. The temperature of the CFT is equal to the Hawking temperature of the black hole. The total energy of the CFT system $\bar{E}_{\beta}$ is equal to the mass of the black hole $M=\mu/\tilde{G}_{N}$ Balasubramanian:1999re . From the definition of $\bar{E}=(E_{n}+E_{m})/2$ and $\omega=E_{m}-E_{n}$, $|\omega|\leq 2\bar{E}$. But we do not have access to the discrete energy spectrum of quantized gravity. $\bar{E}$ (energy of heavy operators) is enhanced by a $1/G_{N}$ factor as compared to $\omega$ (energy of light operators) when measured in the unit of $1/R$. Hence, we can effectively take $\omega\in(-\infty,\infty)$. The use of these two different energy scales were implicit in other works, for example, see Horowitz:1999jd ; Jokela:2015sza .
@@
In this work, we will consider perturbation with a light operator on the CFT. We are working with linear response theory so the strength of the source is chosen to be small. It has been shown that strong global perturbation with a light operator can form black holes Bhattacharyya:2009uu . One can use geodesic approximation Balasubramanian:1999zv in case of perturbation with heavy operators. The technical challenge is that there are no real geodesic which connects two timelike boundary points Balasubramanian:2012tu .
@@
We use the spectral function $A(\beta,\omega)$ of a scalar operator in the holographic CFT to calculate $f(\bar{E})$ of the scalar operator. In a thermal state with inverse temperature $\beta$ and energy expectation value $\bar{E}_{\beta}$,
$A(\beta,\omega)=2\sinh(\beta\omega/2)f(\bar{E}_{\beta},\omega)^{2},\qquad A(\beta,\omega)=-2\,\text{Im}\;\tilde{G}_{R}(\beta,\omega)$ (2.5)
$G_{R}(t,t^{\prime})=-i\theta(t-t^{\prime})\langle[\mathcal{O}(t),\mathcal{O}(t^{\prime})]\rangle_{\beta}$ (2.6)
@@
The retarded Green function $\tilde{G}_{R}(\beta,\omega)$ is calculated using the Son-Starinets prescription Son:2002sd . We consider a minimally-coupled massive scalar field $\phi$ of mass squared $m^{2}$ in the $d+1$-D bulk spacetime. The equation of motion in terms of the Fourier modes of $t$ and the spherical harmonics of the (d1)(d-1)-sphere is
@@
$\qquad\tilde{G}_{R,\Delta+}(\beta,\omega)=-\frac{\mathcal{B}(\omega,l)}{\mathcal{A}(\omega,l)}$ (2.12)
For brevity, we have suppressed the $l$-dependence in $\tilde{G}_{R,\Delta+}(\beta,\omega)$, and instead, emphasized $\beta$-dependence. If $\mathcal{O}_{\Delta-}$ is well defined, then its retarded Green function is given by
$\tilde{G}_{R,\Delta-}(\beta,\omega)=\frac{\mathcal{A}(\omega,l)}{\mathcal{B}(\omega,l)}$ (2.13)
The $f$-function can be calculated using (2.5), where we have suppressed the $l$-dependence. We are looking for $\mu$, $\beta$, $T$ or $r_{h}$ dependence. While in other works, $\omega$-dependence is the priority. So it is important that (2.12) is nicely normalized $\text{lim}_{t\to 0+}G_{R}(t)=-i/t^{2\Delta}$, while using the $r$ coordinate Dodelson:2022yvn . It should be noted that a different coordinate system was used in Son:2002sd . For only the spectral function, we can also perform the calculation near the horizon due to flux conservation. In the rest of the paper, we will consider an operator $\mathcal{O}_{\Delta}$ of conformal weight $\Delta$ which can be either $\Delta_{+}$ or $\Delta_{-}$.
We will be varying $\mu$ so we are really working with $f(\bar{E})=f(\mu/\tilde{G}_{N})=\hat{f}(\mu)$. With a slight abuse of notation, we will still use $f(\mu)$ instead of $\hat{f}(\mu)$. The quantity that appears in physical processes is $f^{\prime}(\bar{E})/f(\bar{E})$. So we will use the relation